# Observational properties of black hole in quantum fluctuation modified gravity

### v1

@@ 3.1 The radiant energy flux
The QFMGBH cannot be directly observed, but the physical properties of the accretion disks around it are significant. We can further limit the range of $\alpha$ by the physical properties of the disks, for example the radiant energy flux. In other words, the numerical values of the radiant energy flux should be real numbers, not complex numbers. We find that the limited range of $\alpha$ for $\omega=1/3$ is $(-1.00,0.74]$, for $\omega=0$ is $[0,0.66]$, for $\omega=-2/3$ is $[0.45,0.54]$, and for $\omega=-4/3$ is $[0.47,0.53]$. These limited ranges also fulfill the asymptotically flat condition. Here we note that: $(\textbf{1})$ when $\omega=0$ and $\alpha=0$, the QFMGBH spacetime becomes the Schwarzschild-like spacetime with effective mass $2M_{\rm{eff}}=2M-D$; $(\textbf{2})$ when $\alpha=0.50$, as $\omega$ changes, $P$ always equals $-4$, and the QFMGBH spacetime remains unchanged.
@@ 3.2 The radiation temperature
3.3 The observed luminosity
The luminosity $L(\nu)$ of a thin accretion disk surrounding a BH is inferred from the red-shifted black-body spectrum and can be defined as [64].
[…]
@@ 3.3 The observed luminosity
$1+z=\frac{1+\Omega r\sin\phi\sin\gamma}{\sqrt{-g_{tt}-\Omega^{2}g_{\phi\phi}}}$ (3.4)
The bending effect of light is ignored[65, 66].
@@
3.4 The radiative efficiency
Now, we investigate the radiative efficiency $\epsilon$. During mass accretion around a QFMGBH, the radiative efficiency is a key parameter that characterizes the central object’s ability to convert its rest mass into outward radiating energy. We can define the radiative efficiency as the ratio of two rates: the rate at which radiation energy escapes from the disk surface to infinity, and the rate of mass energy transfer of the central dense object during mass accretion [61, 62]. We assume that all photons can escape to infinity, so the radiative efficiency depends solely on the specific energy of the particle at the innermost stable circular orbit, $r_{\rm{ISCO}}$ [61, 62].
$\epsilon=1-E_{\rm{ISCO}}$ (3.5)
@@ 3.4 The radiative efficiency
3.5 Some important data
@@
GR -- $6.00$ $1.37\times 10^{-5}$ $0.095$ $0.033$ $0.0086$ $0.057$
$1/3$ $-0.8$ $3.09$ $8.01\times 10^{-5}$ $0.148$ $0.048$ $0.0139$ $0.091$
$-0.4$ $3.33$ $6.83\times 10^{-5}$ $0.143$ $0.046$ $0.0137$ $0.089$
@@
3.6 The shadow
In this section, we focus on the shadow images of QFMGBHs. The BH is surrounded by various objects (e.g., the accretion disk) that emit photons, which can be received by an observer at infinity, providing information about the BH and the surrounding matter. To obtain shadow images of QFMGBHs, we use an open-source general relativistic ray-tracing code called GYOTO [25, 67]. The original GYOTO code for the thin accretion disk, which is geometrically thin but optically thick, describes the Kerr spacetime. We modify the original code to simulate images of the geometrically thin and optically thick accretion disk around QFMGBHs.
@@ 3.6 The shadow
In figures 10, 11, 12, and 13, we present the accretion disks obtained using GYOTO, with the observation radius set to $r_{\rm o}=80M$, the observation angle $\theta_{\rm o}=70^{\circ}$. The accretion disks in figures 10, 11, and 12 correspond to fixed values of $\omega=1/3$, $\omega=0$, and $\omega=-2/3$, respectively. From figures 10, 11, and 12, we observe that as $\alpha$ increases, the area of the shadow of the BH increases. Figure 13 corresponds to a fixed value of $\omega=-4/3$, and from which, we observe that as $\alpha$ increases, the area of the shadow of the QFMGBH decreases. Furthermore, by comparing figures 10, 11, 12, and 13, we find that as $\omega$ decreases, the absolute value of the rate of change of the area increases. Notably, for all four values of $\omega$, the area of the shadow of the QFMGBH is smaller than that of the Schwarzschild BH.
@@ 4 Conclusions
Acknowledgments
This study is supported in part by National Natural Science Foundation of China (Grant No. 12333008) and Hebei Provincial Natural Science Foundation of China (Grant No. A2021201034).
Data Availability Statement: No Data associated in the manuscript.

### v2

@@ III.1 The radiant energy flux
The QFMGBH cannot be directly observed, but the physical properties of the accretion disks around it are significant. We can further limit the range of $\alpha$ by the physical properties of the disks, for example the radiant energy flux. In other words, the numerical values of the radiant energy flux should be real numbers, not complex numbers. We find that the limited range of $\alpha$ for $\omega=1/3$ is (1.00,0.74](-1.00,0.74], for $\omega=0$ is $[0,0.66]$, for $\omega=-2/3$ is $[0.45,0.54]$, and for $\omega=-4/3$ is $[0.47,0.53]$. These limited ranges also fulfill the asymptotically flat condition. Here we note that: $(\textbf{1})$ when $\omega=0$ and $\alpha=0$, the QFMGBH spacetime becomes the Schwarzschild-like spacetime with effective mass $2M_{\rm{eff}}=2M-D$; $(\textbf{2})$ when $\alpha=0.50$, as $\omega$ changes, $P$ always equals 4-4, and the QFMGBH spacetime remains unchanged.
@@
From these discussions, we see that there is degeneracy between parameters $\omega$ and $\alpha$. The general approach to reducing degeneracies between parameters is to fix some parameters while examining the impacts of changes in other parameters, as we have done in the analysis of the graph. In Chang:2024cbi , Bayesian analysis was adopted to constrain the parameters based on the analytic dual-cone accretion model, however, some parameters were also fixed at the same time. In vanderGucht:2019bhl ; Fang:2024hbw ; Janssen:2025rbl ; Sharipov:2025dgh , machine learning methods were used to limit parameters, and generally some parameters were needed to be fixed in order to break the degeneracy between parameters, see for example vanderGucht:2019bhl .
@@ III.2 The radiation temperature
III.3 The observed luminosity
The luminosity $L(\nu)$ of a thin accretion disk surrounding a BH is inferred from the red-shifted black-body spectrum and can be defined as Torres_2002 .
$L(\nu)=4\pi d^{2}I(\nu)=\frac{8\pi h\cos\gamma}{c^{2}}\int_{r_{\rm i}}^{r_{\rm f}}\int_{0}^{2\pi}\frac{\nu_{\rm e}^{3}r}{e^{\frac{h\nu_{\rm e}}{k_{\rm B}T}}-1}\,d\phi\,dr$ (20)
@@ III.3 The observed luminosity
$1+z=\frac{1+\Omega r\sin\phi\sin\gamma}{\sqrt{-g_{tt}-\Omega^{2}g_{\phi\phi}}}$ (21)
The bending effect of light is ignoredarticle ; Bhattacharyya_2001 .
@@
From the above discussion, it can be inferred that the energy flux, the temperature, and the emission spectrum are higher from the QFMGBH as compared to the Schwarzschild BH. This is because that the $r_{\rm{ISCO}}$ of QFMGBH is smaller than or equal to that of Schwarzschild BH, so the disk around a QFMGBH is closer to the BH.
@@
III.4 The radiative efficiency
Now, we investigate the radiative efficiency $\epsilon$. During mass accretion around a QFMGBH, the radiative efficiency is a key parameter that characterizes the central object’s ability to convert its rest mass into outward radiating energy. We can define the radiative efficiency as the ratio of two rates: the rate at which radiation energy escapes from the disk surface to infinity, and the rate of mass energy transfer of the central dense object during mass accretion Novikov:1973kta ; Page:1974he . We assume that all photons can escape to infinity, so the radiative efficiency depends solely on the specific energy of the particle at the innermost stable circular orbit, $r_{\rm{ISCO}}$ Novikov:1973kta ; Page:1974he .
$\epsilon=1-E_{\rm{ISCO}}$ (22)
@@ III.4 The radiative efficiency
III.5 Some important data
@@
GR - $6.00$ $1.37\times 10^{-5}$ $0.095$ $0.033$ $0.0086$ $0.057$
$1/3$ 0.8-0.8 $3.09$ $8.01\times 10^{-5}$ $0.148$ $0.048$ $0.0139$ $0.091$
0.4-0.4 $3.33$ $6.83\times 10^{-5}$ $0.143$ $0.046$ $0.0137$ $0.089$
@@
III.6 The shadow
In this section, we focus on the shadow images of QFMGBHs. The BH is surrounded by various objects (e.g., the accretion disk) that emit photons, which can be received by an observer at infinity, providing information about the BH and the surrounding matter. To obtain shadow images of QFMGBHs, we use an open-source general relativistic ray-tracing code called GYOTO Vincent_2011 ; aimar2023gyoto20polarizedrelativistic . The original GYOTO code for the thin accretion disk, which is geometrically thin but optically thick, describes the Kerr spacetime. We modify the original code to simulate images of the geometrically thin and optically thick accretion disk around QFMGBHs.
@@ III.6 The shadow
In figures 10, 11, 12, and 13, we present the accretion disks obtained using GYOTO, with the observation radius set to $r_{\rm o}=80M$, the observation angle $\theta_{\rm o}=70^{\circ}$. The accretion disks in figures 10, 11, and 12 correspond to fixed values of $\omega=1/3$, $\omega=0$, and $\omega=-2/3$, respectively. For comparison, we also present the shadow of the Schwarzschild BH (no fluctuations and no fluids around it). From figures 10, 11, and 12, we observe that as $\alpha$ increases, the area of the shadow of the BH increases. Figure 11 with $\omega=0$ show the shadow of Schwarzschild like BH with reduced mass 2MD2M-D. Figure 13 corresponds to a fixed value of $\omega=-4/3$, and from which, we observe that as $\alpha$ increases, the area of the shadow of the QFMGBH decreases. Furthermore, by comparing figures 10, 11, 12, and 13, we find that as $\omega$ decreases, the absolute value of the rate of change of the area increases. Notably, for all four values of $\omega$, the area of the shadow of the QFMGBH is smaller than that of the Schwarzschild BH.
@@ IV Conclusions
Acknowledgements.
This study is supported in part by National Natural Science Foundation of China (Grant No. 12333008) and Hebei Provincial Natural Science Foundation of China (Grant No. A2021201034). The GYOTO code can requested from us by email.